OpenStudy (anonymous):

A fruit display in a market is in the shape of a pyramid with one peice of fruit at the top. Each layer below the top forms a square shape with one additional piece of fruit per side. If a total of 55 pieces of fruit are used , and a single piece of fruit at the top counts as a layer,how many layers of fruit will there be? PLEASE SHOW WORK :]

OpenStudy (carlsmith):

Homework?

OpenStudy (anonymous):

no just studying this is a question on the practice test can u figure it out i really need help..and please show work ;]

OpenStudy (carlsmith):

This isn't really so hard to figure out. Each layer contains the number of apples on one side times by itself, or squared. There's one on the top layer, 1 * 1 = 1 ~ two on the next layer, 2 * 2 = 4 and three in the next layer, 3 * 3 = 9 and so on. 1, 4, 9, 16, 25. Just keep adding each layer's apples together until you run out of apples.

OpenStudy (anonymous):

Layer 1 - 1 Layer 2 - 4 Layer 3 - 9 Layer 4 - 16 Layer 5 - 25 Total = 55, so 5 layers!

OpenStudy (anonymous):

thank you i just thought the question was worded weirdly

OpenStudy (carlsmith):

No problem. All the best.